# A Role of Environmental Complexity on Representation Learning in Deep Reinforcement Learning Agents

### v1

@@ Abstract
The environments where individuals live can present diverse navigation challenges, resulting in varying navigation abilities and strategies. Inspired by differing urban layouts and the Dual Solutions Paradigm test used for human navigators, we developed a simulated navigation environment to train deep reinforcement learning agents in a shortcut usage task. We modulated the frequency of exposure to a shortcut and navigation cue, leading to the development of artificial agents with differing abilities. We examined the encoded representations in artificial neural networks driving these agents, revealing intricate dynamics in representation learning, and correlated them with shortcut use preferences. Furthermore, we demonstrated methods to analyze representations across a population of nodes, which proved effective in finding patterns in what would otherwise be noisy single-node data. These techniques may also have broader applications in studying neural activity. From our observations in representation learning dynamics, we propose insights for human navigation learning, emphasizing the importance of navigation challenges in developing strong landmark knowledge over repeated exposures to landmarks alone.
@@ 1 Introduction
There are many factors that can contribute to individual differences in navigation skill, with sex [19] and age [15] being well-studied aspects. The primary difference we focus on in this study is that of an individual’s home environment, which can also have profound effects on navigation ability. Work from Barhorst-Cates et al. [1] explored differences in skill between participants from Salt Lake City (Utah, USA) with its regular, grid-like streets and participants from Padua (Veneto, Italy) with organic, maze-like streets. They generally found that individuals from Padua had better access to survey knowledge and better ability to use cues. The primary experimental protocol performed with humans we focus on is known as the Dual Solutions Paradigm (DSP) [12]. This task evaluates survey knowledge by first guiding participants on a standard route along the outside of a maze, and later assessing how often they attempt to find novel paths (shortcuts) between locations.
In this study, we apply deep reinforcement learning (RL) [23] to model the navigation learning process. Our dual objectives are to leverage human navigation understanding to construct a training environment for artificial agents, and to develop tools for probing the representations that the agents learn. RL has been successfully applied to various complex navigation tasks, providing a robust challenge for testing agent capabilities [6, 16, 17]. Inspired by the DSP, we developed a simulated environment to train RL agents, modulating shortcut availability to abstractly reflect differences in environment complexity seen in Salt Lake City and Padua. We examined how variations in the “growing up” setting of an agent influenced learning dynamics and strategy development. Specifically, we focused on shortcut usage as a measure of learning and behavior, and drew correlations between performance and representation learning.
In past work, we successfully developed methods to analyze the neural network representations developed in navigating agents [9], and we continue building more analysis and exploration tools in the present work. Representations can be viewed as correlations between the state of the environment and the activity of the agent’s neural network. In other words, they act as an internal model of the world that the agent faces. We explored spatial representations recorded from individual nodes, inspired by place cells found in the hippocampal system [18]. We also examined landmark sensitive representations, motivated by the importance of landmark usage in human navigation. Beyond individual neural network nodes, we also looked at two representations found across populations of nodes. We find that looking at the activity of a population smooths over noise that individual node activities have, and can even lend context to better understand what those noisy activations encode. We suggest that similar techniques to ours may be useful in general to understand the activities of neural networks as well as neuronal firing rates in brains.
Our results show that different representations evolve at various training stages, with environmental complexity affecting learning dynamics nonlinearly. First, basic spatial maps in individual nodes (reminiscent of hippocampal place cells) develop early in training, while landmark sensitivity develops later, and a trajectory sensitive representation found at the population level continues to develop even after the RL algorithm has converged to a near-optimal policy. We also observe complex learning dynamics in landmark sensitivity. We find that frequent presentation of a salient cue leads to a stable encoding of the cue. However, if another agent is exposed to the cue less often but in a context of greater navigation difficulty, this stimulates the network to eventually learn to encode the cue more robustly than frequent presentation alone.
Building on these insights from artificial agents, we make extrapolations about human navigators from environments of different navigation difficulty levels performing the DSP task. We predict that individuals from environments with consistently available distal landmarks, like Salt Lake City, will have better landmark knowledge compared to counterparts from more complex environments like Padua when both groups have limited navigation experience. On the other hand, when looking at groups with more experience or higher skill levels, those from Padua, which presents greater navigational challenge, will have stronger landmark knowledge.
Lastly, we observed significant changes in learning efficiency in artificial agents triggered by small changes to the environment. This observation leads us to propose a modification to the DSP where a few demonstrations of shortcuts are given initially. We hypothesize that this adjustment would strongly boost the frequency of shortcut usage in those who might otherwise be route followers.
@@ 2.1 Reinforcement Learning
In this paper we train artificial agents with deep RL [23] to perform a navigation task. The overall set up and training approach are similar to those introduced in an earlier publication [9]. The agents are trained in an environment that can be defined as a Partially Observable Markov Decision Process (POMDP) defined by a $(S,A,P,R,\Omega,O)$ tuple. At each time step $t$, the environment has state $s_{t}\in S$ and gives a limited amount of information about that state to the agent in an observation $o_{t}\in\Omega$. The mapping $O:S\rightarrow\Omega$ defines what information is given. The agent performs actions from a discrete action space $a_{t}\in A$, the state evolves according to a transition function $P:S\times A\rightarrow S$, and a reward $r_{t}$ is given according to $R:S\times A\rightarrow\mathbb{R}$.
@@ 2.2 The simulated navigation environment
In this study, we train agents to perform a 2D navigation task in Python, depicted in Fig 1A and 1B. The agent’s goal is to navigate to a target located in the top-right corner of the maze, represented as a gray box in the drawing. Sight lines (visualized in Fig 1) return both the color and distance to the wall each line intersects. This information forms a 24-dimensional observation vector $o_{t}\in\Omega=\mathbb{R}^{24}$, as each of the 12 vision lines provides two scalar values. At each time step, the agent picks from four possible actions $a_{t}$: a left or right turn of $0.2$ radians, a forward movement of $10$ units, or no movement. The navigation arena is a $[0,300]\times[0,300]$ square.
The task is structured as an episodic RL task [23]. Each episode resets either when the agent reaches the goal or when $200$ time steps elapse. If the goal is reached, a reward $r_{t}$ of $1$ is earned, and the reward is 00 for all other time steps. Upon resetting, the starting position of the agent is randomized such that the starting height is in $[0,240]$ to ensure that it starts below the corridor, which is the area above the maze’s long wall, 50 units high from the top edge. An initial direction is also selected uniformly at random in $[0,2\pi)$.
This environment comes with a key parameter $p$, which determines the probability of the shortcut being open in any given episode. This parameter gives variations of the environment, and each agent is trained with a fixed $p$ throughout all episodes. The state of the environment $s_{t}$ is comprised of the agent’s location, direction, and whether the shortcut is open or not. After any initial randomizations of the state, the state evolves according to a deterministic transition function $P$ that updates the agent’s location and heading based on the action $a_{t}$ that it picks.
Figure 1: A and B. Depictions of the simulated shortcut navigation environment with A. shortcut closed and B. shortcut opened. All walls are colored white except for the shortcut wall when closed, which is purple. A gray box in the very top-right corner represents the navigation target. The agent is depicted as a yellow triangle, with white lines extending outwards its sight lines. C. A schematic of the neural network that agents are parameterized by. Each block represents network layer of 64 nodes, and the arrows represent the activations of one layer being passed to the next in a fully-connected weighted sum. The block with a circled arrow is a recurrent layer. Observation are input ($o_{t}$ on the left of the schematic), and the network splits into actor ($\pi$) and critic ($V$) branches.
@@ 2.4 Learning curves and initial non-learning periods
Learning curves, which graphically represent the performance of agents based on their cumulative experienced time steps, are shown in Fig 2A. These show mean episode length for individual agents, where an episode length of 200 means the time limit was reached before the agent made it to the target. At the start of training, agents perform random actions due to their randomly initialized neural networks. The first meaningful learning signals occur when agents accidentally reach the goal, from which they quickly learn to replicate successful actions. The learning curves show a steep increase in performance as the agents develop effective strategies at this point.
This pattern leads to an initial non-performant period which has random length, and its length does not impact learning dynamics afterwards. Since the shortcut path is shorter, it is easier for an agent taking random actions to reach the platform when the shortcut is open. Hence, agents trained with a higher probability $p$ of the shortcut being open typically experience a shorter non-performant period. We disregard this initial variability by looking for the first time an agent’s average episode length drops below 180 time steps, shown by dots in Fig 2A. Fig 2B adjusts learning curves to start from this point, providing a clearer view of learning progression. Each agent undergoes training for $66e$ 6 time steps beyond this initial point, and all results in this paper will use these shifted starting points.
@@ 3.1 Training conditions affect navigation ability
Figure 3: (A-C) Performance curves for agents trained with different $p$ probabilities. A. Mean episode length on closed shortcut episodes. B. Mean episode length on open shortcut episodes. C. Mean shortcut use rate on open shortcut episodes. D. Shortcut use rate compared to episode length on open shortcut episodes, with each point representing one agent from one checkpoint.
In this section, we begin by assessing how efficiently agents are able to learn the shortcut navigation task, with evaluation procedures detailed in the Methods section. Performance on closed and open shortcut episodes are shown in Fig 3A and Fig 3B respectively. A few key observations can be made from these plots. First, Fig 3A demonstrates that regardless of the value of $p$, agents rapidly achieve optimal performance in closed shortcut trials once initial learning signs appear. Agents with $p=0.1$ agents appear to reach an optimal policy slightly faster than others, likely due to more frequent exposure to closed shortcut episode. However, learning speeds in this condition for agents with $p\geq 0.2$ are all roughly equivalent.
Fig 3B and C show navigation speed and shortcut use rate respectively on open shortcut episodes. Fig 3D confirms a clear correlation between the two. Agents trained in higher $p$ environments gain more experience with open shortcut trials, directly influencing the speed at which they learn to use the shortcut consistently. Given enough training time, agents can learn to use shortcuts at a rate greater than the probability that they saw shortcuts during training. For instance, $p=0.1$ agents achieve around $50$% shortcut usage by the end of training, and agents trained with $p\geq 0.3$ all reach nearly $100$% shortcut usage. Given an even longer training duration, it is possible that all agents could converge to fully optimal policies.
@@ 3.2 Position Representations
Representations can be thought of as encodings of environment state information within the activations of an artificial neural network layer [2, 10]. Using this definition, we may say that the neural network represents a d-dimensional vector feature $x\in\mathbb{R}^{d}$ if there is a mapping $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ that maps the $n$ activations of a network layer to that feature. In practice, finding a formal mapping is impractical, so instead we focus on measuring the correlation between network activity and state or environmental information. Sometimes, we call the information about state a feature of the environment. When a strong correlation exists we consider the network activity to represent that feature.
Drawing inspiration from the hippocampal place cells and grid cells in the brain, which activate selectively based on spatial location [18], we initially explore position-based encodings in our artificial agents. These nodes help form a spatial representation of an agent’s position. We start by detailing our approach to computing these representations and then develop a measure of spatial representation strength. We discuss how having spatially consistent activations is necessary for basic navigation, but insufficient for advanced strategies. Then, we describe a technique that examines a population of nodes, giving us clearer insight into the spatial features that artificial agents tend to encode. This approach helps uncover representations which could otherwise be difficult to identify when looking at noisy activation patterns in individual nodes.
For all discussions of representations, we analyze activities in the recurrent layer of an agent’s neural network. This is the last layer in sequence that is shared between the actor and critic sides of the network. In principle, representations could be searched for in any layer of the neural network, and other works have performed representation analysis on for example, the penultimate actor layer [2]. Our focus remains on the recurrent layer due to its role in integrating and storing information, acting as the agent’s internal memory.
@@ 3.2.1 Spatial heatmaps and local variance measure
The first individual node level representations we present are spatial representations, examined by generating spatial activation heatmaps. While tasking agents with completing the 100 evaluation episodes outlined in Sec 3.1, we collect both the positions that agents were in and activations of their recurrent neural network layers at each time step. Heatmaps are generated by taking a Gaussian-weighted spatial average of these individual activation data points. We partition the maze space into a 30 x 30 grid, coloring each grid point by a spatially-weighted average of activations from a single node across all episodes. In other words, a heatmap illustrates how a node tended to activate when the agent was in different parts of the environment. Heatmaps are standardized to have zero mean and unit variance, highlighting deviations from a node’s typical outputs rather than absolute activation values. This gives us a clearer idea of what the node encodes.
Examples of some heatmaps are shown in Fig 4. These are chosen to be representative of common types of heatmaps observed in agents, but without preference for the $p$ environment or point in training used. Some nodes show clear sensitivities for certain parts of the maze, such as corners (Fig 4A), the corridor (Fig 4B, E, G), nearby the corridor entrance (Fig 4C, D) or have a center-surround configuration (Fig 4F). They also range in amounts of noisiness, from what we might consider easy to interpret (Fig 4A-F) to noisy and difficult to interpret (Fig 4G-H).
@@
To quantify the spatial representation strength of a node, we introduce a measure called the mean local variance. This measures the local consistency of activation patterns within the heatmap. As illustrated in Fig 5A, we calculate standard deviations for each 5x5 pixel window in the heatmap and average these to derive a single score for each node’s spatial representation. The score effectively captures the intuition that nodes with more consistent and interpretable spatial activations more strongly encode location information. It is likely that a single node is involved in the computation of multiple representations simultaneously, and a noisier heatmap may indicate a node that does not encode spatial information as much as other features. Note that while this score uses the standard deviation, we call it a local variance, but the two terms are equivalent in interpretation.
@@
Fig 6A presents how the mean local variance measure evolves over training for $p=0.1$ agents, alongside the $p=0.1$ closed episode length performance curve from Fig 3A. We observe a sharp decline in both measures early in training as agents become proficient at navigation, followed by a plateau, suggesting no further improvement. A clear correlation between these two is shown in Fig 6B, and holds across all $p$ agents. One may think of the closed shortcut performance curves as a proxy for basic navigation skill development since there is only one path to take, and navigation is relatively easy on these episodes. Spatial representation strength develops simultaneously with closed shortcut performance, suggesting that stable location encodings are crucial for basic navigation. We might similarly expect that in animals and humans, spatial consistency develops early in training, setting the foundation for other types of representations that mature more gradually.

### v2

@@ Abstract
We developed a simulated environment to train deep reinforcement learning agents on a shortcut usage navigation task, motivated by the Dual Solutions Paradigm test used for human navigators. We manipulated the frequency with which agents were exposed to a shortcut and a navigation cue, to investigate how these factors influence shortcut usage development. We find that all agents rapidly achieve optimal performance in closed shortcut trials once initial learning starts. However, their navigation speed and shortcut usage when it is open happen faster in agents with higher shortcut exposure. Analysis of the agents’ artificial neural networks activity revealed that frequent presentation of a cue initially resulted in better encoding of the cue in the activity of individual nodes, compared to agents who encountered the cue less often. However, stronger cue representations were ultimately formed through the use of the cue in the context of navigation planning, rather than simply through exposure. We found that in all agents, spatial representations develop early in training and subsequently stabilize before navigation strategies fully develop, suggesting that having spatially consistent activations is necessary for basic navigation, but insufficient for advanced strategies. Further, using new analysis techniques, we found that the planned trajectory rather than the agent’s immediate location is encoded in the agent’s networks. Moreover, the encoding is represented at the population rather than the individual node level. These techniques could have broader applications in studying neural activity across populations of neurons or network nodes beyond individual activity patterns.
@@ 1 Introduction
There are many factors that can contribute to individual differences in navigation skill, including, e.g., sex [19] and age [15]. Recently it has been suggested [1] that individual’s home environment also helps to shape their navigation strategies and skills. Barhorst-Cates et al. [1] compared navigation performance between people from Salt Lake City (Utah, USA) and those from Padua (Veneto, Italy). These two cities provide contrasting environments (sample maps are shown in Fig 1): Salt Lake City has grid-like streets and surrounding mountains provide global cues and Padua has organic, maze-like streets. It was found that individuals from Padua generally had better access to survey knowledge, ability to use cues, and overall navigation skill. One of the tasks that was used in this study is the dual-solutions paradigm (DSP) [12], which assesses how often shortcuts are used during navigation in a maze. In DSP the tendency to use shortcuts is thought to indicate the use of survey knowledge by the individual, and it has been used in both rodent and human studies.
Inspired by the human studies, we developed a toy version of these navigation tasks to study the effect of the “home” environment, the shortcut usage, and the attendant “survey knowledge” in the artificial agents (in Section 2.2). We applied deep reinforcement learning (RL) [23] to model the navigation learning process in artificial agents. The agents were trained with different availability of a visual landmark and a shortcut, resembling different home environments for humans, and then were tested in a toy version of the dual-solution paradigm task. We examined their navigation strategies and shortcut use, and analyzed the resulting activities in their artificial neural networks for representations of survey knowledge and other components of navigation decisions.
Figure 1: Maps of A. Padua (Veneto, Italy) and B. Salt Lake City (Utah, USA), taken from Google Maps [5]
Our dual objectives are to leverage understanding of human navigation to construct an informative and constrained training environment for artificial agents, and to develop tools for probing and analyzing the network activities (representations) at different stages of the agents’ training.
RL has been successfully applied to various complex navigation tasks, providing a robust challenge for testing agent capabilities [6, 16, 17]. In our approach [9] we use simplified versions of the navigating environments, to enable a more nuanced understanding of the learning process, the resulting strategies, and the analysis of the neural network representations in navigating agents [9].
Representations can be viewed as correlations between the state of the environment and the activity of the agent’s neural network. In other words, they act as an internal model of the world that the agent faces. In this work, we first looked at the spatial representations recorded from individual nodes. There are similar to activities of place cells found in the hippocampal system [18]. We find that these encodings are foundational to any navigation strategy and develop early in training. In contrast, landmark encoding, which is measured as activity in nodes that differentiate between the presence and absence of a given visual cue, continues to develop throughout the training period. In particular, the landmark encoding is stronger in agents who are given more access to landmarks early on, but continue to develop more robustly in agents that learn to apply them as part of their navigation strategy, independent of the frequency that agents encounter the cues.
Beyond individual neural network nodes, we also looked at two representations coded across populations of nodes. First, we show that clustering the collection of spatial activity heatmaps from all network nodes can be used to reveal what salient features of the environment are being encoded by the network. Individual node activity patterns can be noisy and hard-to-interpret. Our clustering analysis allows to look beyond the noise to reveal the features of the environment that are available to the agent. This analysis can serve as a useful exploration tool to understand activity in brains or artificial neural networks of agents. Second, we find that agents are able to effectively encode the trajectory that they are on, suggesting an intent or planned path, using a population representation. This planned path representation is continuously refined throughout the training process. It is also independent of cue availability, which we can show by making use of our ability to manipulate the visual cues given to agents during navigation.
While this experimental set up is much abstracted from the more complex navigation world that real people and animals face, the results may still provide some insight into biological navigators. Our observations on landmark sensitive nodes suggest suggest that individuals from Padua who may have less consistent landmarks than those in Salt Lake City nevertheless can develop stronger internal cue representations due to their increased survey knowledge of experiencing difficult navigating conditions. From the variations of our toy model, we observed significant changes in learning efficiency in artificial agents triggered by relatively small changes to the environment. We propose a modification to the DSP where a few demonstrations of shortcuts are given initially and hypothesize that this adjustment would strongly boost the frequency of shortcut usage in those subjects who might otherwise be route followers.
Overall, we find that even relatively simple navigation environments can give rise to interesting behavior and representation learning dynamics in artificial agents. These trained neural networks can provide insight into the internal cognition of real navigators, and suggest interventions to affect skill and knowledge development.
@@ 2.1 Reinforcement Learning
We train deep RL [23] artificial agents to perform a navigation task, using a set up and training approach similar to ones we used in an earlier publication [9]. The agents are trained in an environment that can be conceptualized as a Partially Observable Markov Decision Process (POMDP) defined by a $(S,A,P,R,\Omega,O)$ tuple. At each time step $t$, the environment has state $s_{t}\in S$ and gives a limited amount of information about that state to the agent in an observation $o_{t}\in\Omega$. The mapping $O:S\rightarrow\Omega$ defines what information is given. The agent performs actions from a discrete action space $a_{t}\in A$, the state evolves according to a transition function $P:S\times A\rightarrow S$, and a reward $r_{t}$ is given according to $R:S\times A\rightarrow\mathbb{R}$.
@@ 2.2 The simulated navigation environment
In this study, we train agents to perform an episodic [23] navigation task (simulations run in Python), depicted in Fig 2A and 2B. The agent’s goal is to navigate to an invisible target located in the top-right corner of the maze, represented as a gray box in the drawing, at the end of a corridor at the top of the environment. A key feature of this environment is the pink wall in the center of the corridor (Fig 2A). This wall can be opened (Fig 2B), creating a shortcut in the maze and simultaneously removing the primary salient landmark in the environment. At the start of the episode, the shortcut is randomly chosen to be open with probability $p$, where $p$ serves as the key parameter of the environment. Varying this parameter gives rise to variations of the environment, and each agent is trained with a fixed $p$ throughout all episodes.
Sight lines (visualized in Fig 2) return both the color and distance to the wall each line intersects. Color is encoded as a one-hot vector with two possibilities (white or pink). This information forms a 36-dimensional observation vector $o_{t}\in\Omega=\mathbb{R}^{36}$, as each of the 12 vision lines provides one scalar value (distance) and two dimensions for color. At each time step, the agent picks from four possible actions $a_{t}$: a left or right turn of $0.2$ radians, a forward movement of $10$ units, or no movement. The navigation arena is a $[0,300]\times[0,300]$ square.
Each episode resets either when the agent reaches the goal or when $200$ time steps elapse. If the goal is reached, a reward $r_{t}$ of $1$ is earned, and the reward is 00 for all other time steps. Upon resetting, the starting position of the agent is randomized such that the starting height is in $[0,240]$ to ensure that it starts below the corridor, which is the area above the maze’s long wall, 50 units high from the top edge. An initial direction is also selected uniformly at random in $[0,2\pi)$. After initialization, the environment evolves deterministically according to actions selected by the agent.
Figure 2: A and B. Depictions of the simulated shortcut navigation environment with A. shortcut closed and B. shortcut opened. All walls are colored white except for the shortcut wall when closed, which is pink. A gray box in the very top-right corner represents the navigation target. The agent is depicted as a yellow triangle, with white lines extending outwards its sight lines. C. A schematic of the neural network that agents are parameterized by. Each block represents network layer of 64 nodes, and the arrows represent the activations of one layer being passed to the next in a fully-connected weighted sum. The block with a circled arrow is a recurrent layer. Observation are input ($o_{t}$ on the left of the schematic), and the network splits into actor ($\pi$) and critic ($V$) branches.
@@ 2.4 Learning curves and initial non-learning periods
Learning curves graphically represent the performance of agents based on their cumulative experienced time steps, shown in Fig 3A. These show mean episode length for individual agents, where an episode length of 200 indicates that the time limit was reached and the agent failed to reach the target. At the start of training, agents perform random actions due to their randomly initialized neural networks. The first meaningful learning signals occur when agents first reach the goal, from which they quickly learn to replicate successful trajectories. The learning curves show a steep improvement in performance as the agents develop effective strategies.
This pattern leads to an initial non-performant period which has random length, and its length does not impact learning dynamics afterwards. It is easier for an agent taking random actions to reach the platform when the shortcut is open. Hence, agents trained with a higher probability $p$ of the shortcut being open typically experience a shorter non-performant period. We disregard this initial variability by looking for the first time an agent’s average episode length drops below 180 time steps, shown by dots in Fig 3A. Fig 3B adjusts learning curves to start from this point, providing a clearer view of learning progression. Each agent undergoes training for $66e$ 6 time steps beyond this initial point, and all results in this paper will use these shifted starting points.
@@ 2.5 Evaluation episodes
2.6 Representations
Representations can be thought of as encodings of environment state information within the activations of an artificial neural network layer [2, 10]. Formally, we may say that the neural network represents a d-dimensional vector feature $x\in\mathbb{R}^{d}$ if there is a mapping $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ that maps the $n$ activations of a network layer to that feature. In practice, finding a formal mapping is impractical, so instead we turn to other methods of measuring the information stored in network activity. We measure how strongly individual nodes encode location information based on how consistently they tend to activate in spatial regions. Landmark sensitivity is computed based on the differences in node activation with the pink wall landmark being present or absent. We also examine representations on the population level. Clustering visualizations are used to qualitatively examine what spatial features are being encoded. Finally, we measure distances in population activity space to determine how different trajectories are encoded in networks even if they pass through the same spatial locations. Specific details on how each of these representation measures is calculated is discussed in respective section of the results.
For all discussions of representations, we analyze activities in the recurrent layer of an agent’s neural network. This is the last layer in sequence that is shared between the actor and critic sides of the network. In principle, representations could be searched for in any layer of the neural network, and other works have performed representation analysis on for example, the penultimate actor layer [2]. Our focus remains on the recurrent layer due to its role in integrating and storing information, acting as the agent’s internal memory, and being shared by the actor and critic branches of the network. These properties make it likely to encode attributes generally useful to both policy and value.
@@ 3.1 Training conditions affect navigation ability
Figure 4: (A-C) Performance curves for agents trained with different probabilities of encountering an open shortcut ($p$). A. Mean episode length on closed shortcut episodes. B. Mean episode length on open shortcut episodes. C. Mean shortcut use rate on open shortcut episodes. D. Shortcut use rate compared to episode length on open shortcut episodes, with each point representing one agent from one checkpoint.
We begin by assessing how efficiently agents are able to learn the shortcut navigation task, with evaluation procedures detailed in the Methods section. Performance on closed and open shortcut episodes are shown in Fig 4A and Fig 4B respectively. A few key observations can be made from these plots. First, Fig 4A demonstrates that regardless of frequency of encountering the open shortcut during training (the value of $p$), agents rapidly achieve optimal performance in closed shortcut trials once initial learning starts. Agents trained with $p=0.1$ appear to reach an optimal policy slightly faster than others, due to more frequent exposure to closed shortcut episode. However, learning speeds in this condition for agents with $p\geq 0.2$ are all roughly equivalent.
Fig 4B and C show navigation speed and shortcut use rate respectively on open shortcut episodes. Fig 4D confirms correlation between the two. Agents trained in higher $p$ environments gain more experience with open shortcut trials, directly influencing the speed at which they learn to use the shortcut consistently. Given enough training time, agents can learn to use shortcuts at a rate greater than the probability that they saw shortcuts during training. For instance, $p=0.1$ agents achieve around $50$% shortcut usage by the end of training, and agents trained with $p\geq 0.3$ all reach nearly $100$% shortcut usage. Given even longer training duration, it is possible that all agents could converge to fully optimal policies.
@@ 3.2 Position Representations
Drawing inspiration from the hippocampal place cells and grid cells in the brain, which activate selectively based on spatial location [18], we initially explore position-based encoding in our artificial agents. We start by detailing our approach to computing these representations and then develop a measure of spatial representation strength. Spatial representations develop early in training and subsequently stabilize before navigation strategies fully develop, suggesting that having spatially consistent activations is necessary for basic navigation, but insufficient for advanced strategies. Then, we use clustering methods on spatial activity heatmaps to qualitatively examine the features used by artificial agents in the shortcut task. Clustering methods help uncover features in representations that could otherwise be challenging to identify when looking at noisy activation patterns in individual nodes.
@@ 3.2.1 Spatial heatmaps and local variance measure
We first study spatial representations at the individual-node level by generating spatial activation heatmaps. While an agent is completing the 100 evaluation episodes outlined in Sec 3.1, we collect both the agent’s position and the activations of each node in the recurrent neural network layer at each time step. We first partition the maze space into a 30 x 30 grid and call the center of each grid cell the grid point $\{g_{i}\}$. Then, to generate the spatial heatmap for a node, we take averages of the node’s activities across all collected time steps weighted by proximity between the grid point and the agent at that time. More specifically, let $\{x_{t}\}$ and $\{h_{t,i}\}$ be, respectively, the position of the agent and activity of node $i$ at time $t$. Each grid point in the heatmap is colored based on a Gaussian spatial average of activity. Let $w_{t,i}=\exp(-d^{2}(x_{t},g_{i})/2\sigma^{2})$ where $d$ is Euclidean distance, then the grid point is colored as a weighted average $\bar{h_{g_{i}}}=\sum_{t}h_{t,i}w_{t,i}/\sum_{t}w_{t,i}$. In other words, a heatmap illustrates how a node tended to activate when the agent was in different parts of the environment. Heatmaps are then standardized to have zero mean and unit variance, highlighting deviations from a node’s typical outputs rather than absolute activation values.
Examples of some heatmaps are shown in Fig 5. These are chosen to be representative of common archetypes we observed. Some nodes show clear sensitivities for certain parts of the maze, such as corners (Fig 5A), the corridor (Fig 5B, E, G), nearby the corridor entrance (Fig 5C, D) or have a center-surround configuration (Fig 5F). They also range in amounts of noisiness, from what we might consider easy to interpret (Fig 5A-F) to noisy and difficult to interpret (Fig 5G-H).
@@
To quantify the spatial representation strength of a node, we introduce a measure called the mean local variance. This measures the local consistency of activation patterns within the heatmap. As illustrated in Fig 6A, we calculate standard deviations for each 5x5 pixel window in the heatmap and average these to derive a single score for each node’s spatial representation. The score effectively captures the intuition that nodes with more consistent and interpretable spatial activations (lower mean local variance) more strongly encode location information (see Fig 6B). It is likely that a single node is involved in the computation of multiple representations simultaneously, and a noisier heatmap may indicate a node that does not encode spatial information as much as other features. Note that while this score uses the standard deviation, we call it a local variance, but the two terms are equivalent in interpretation.
@@
Fig 7A presents how the mean local variance measure evolves over training for $p=0.1$ agents, alongside the $p=0.1$ closed episode length performance curve from Fig 4A. We observe a sharp decline in both measures early in training as agents become proficient at navigation, followed by a plateau, suggesting no further improvement. A clear correlation between these two is shown in Fig 7B, and holds across all $p$ agents. One may think of the closed shortcut performance curves as a proxy for basic navigation skill development since there is only one path to take, and navigation is relatively easy on these episodes. Spatial representation strength develops simultaneously with closed shortcut performance, suggesting that stable location encodings are important for basic navigation. In contrast, as we noted in Fig 4C, performance on open shortcut episodes continues to improve long after these encodings stabilize. This suggests that spatial information alone is insufficient for optimal policies in the shortcut environment, and as we will see, other representations are also refined with more training.